ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

19. History of Mathematics

1991 MS Classification: not given Munkholm, Hans J., IMADA, Odense Universitet, Denmark, Munkholm, Ellen S. Poul Heegaard – A home page A home page dedicated to Poul Heegaard’s life and mathematics can be found at

http://www.imada.ou.dk/~hjm/heegaard

Included are, i. a., Heegaard’s unpublished autobiographical notes (in Danish), and an article “Poul Heegaard” to appear in “History of Topology” (I. James, ed.) Elsevier, 1998.